the required margins for your term paper are 3/2 inches at the top and bottom of a paper that has 11 inches of vertical length. how long is the vertical area available for typed information?
Answer by cel072880(4)

the required margins for your term paper are 3/2 inches at the top and bottom of a paper that has 11 inches of vertical length. how long is the vertical area available for typed information?
My Solution:
3/2 + 3/2 = 6/4 total margins at top and bottom
11 - 6/4 = 38/4 or 9 1/2 inches is the total vertical area available for typed information.